OpenStudy (anonymous):

What is the length of stack M N with bar on top ? Round to the nearest tenth of a unit. units

OpenStudy (anonymous):

Graph

OpenStudy (anonymous):

WILL FAN AND MEDAL!!

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

Can you help me @Hero?

hero (hero):

The quickest way to do it is to find the length of the unit segment. Then multiply by the total number of unit segments that span segment MN.

OpenStudy (anonymous):

What??

OpenStudy (anonymous):

@Vocaloid

hero (hero):

The unit segment. If you notice, the entire length of MN can be broken down into six equal segments. Of those six smaller segments, if you take one of them, it would represent the unit segment.

OpenStudy (anonymous):

Can you walk me through using numbers I learn better that way.

hero (hero):

When you have understood that, let me know.

OpenStudy (anonymous):

When I said numbers I meant equations... like @Vocaloid does.

hero (hero):

We'll get to that. But first, do you see that there are six unit segments?

OpenStudy (anonymous):

yes

hero (hero):

Okay, now all we have to do is take just one of those segments create a right triangle and apply pythagorean theorem. Notice the length of the legs of the triangle is just one unit. The pythagorean theorem is a^2 + b^2 = c^2. If we let a and b represent the length of the legs, then \(1^2 + 1^2 = c^2\) or \(1 + 1 = c^2\) or simply \(2 = c^2\). So \(\sqrt{2} = c\)

hero (hero):

\(c\) is the length of one unit segment. Multiply that by 6 and you'll have the length of \(MN\)

OpenStudy (unklerhaukus):

surely it is simpler to apply pythag to the entire interval

OpenStudy (anonymous):

I am SO lost.

hero (hero):

In theory, yes, but if properly applied, this is the simpler method.

hero (hero):

Or the more intuitive method.

OpenStudy (anonymous):

I am NOT good at math at ALL... Sorry.

OpenStudy (unklerhaukus):



OpenStudy (unklerhaukus):